OpenStudy (anonymous):

solve each equation for m, then find the value of m for each value of n, show your work m + 3n = 7; n = -2, 0, 1

OpenStudy (anonymous):

This one is fairly simple. Just do 3xn, -2,0,1- and add whatever is left to get to 7 and that will equal m. Write out 3 separate equations.

OpenStudy (anonymous):

can you show me? the seperate equations.

OpenStudy (anonymous):

3x-2= -6. -6+8=7, m=7. 3x0=0. 0+7=7 m=7 3x1=3. 3+4=7 m=4

OpenStudy (anonymous):

Correction, the first one is m=8, sorry.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Best response? :D

OpenStudy (anonymous):

best response

OpenStudy (anonymous):

wait.... how did you get 8 ? -6 + 8?

OpenStudy (anonymous):

whatever is to the right, after the multiplication, how did you get it?

OpenStudy (anonymous):

Oops! Alright, that one was wrong, sorry. It's 13.

OpenStudy (anonymous):

You are just trying to solve for m in a sense, because n is already given to you 3 times. Just combine the numbers for n, make the equation, and use mental math for the rest. Apparently I wasn't using any, but I made the correction, haha. Hopefully that helped.

OpenStudy (anonymous):

beginning to end man, im even confusing my teacher with this xD

OpenStudy (anonymous):

yeah, thats cool, thanks :P

OpenStudy (anonymous):

Okay so you have, m+3n=7. You would put in one of the numbers it gives you on the right for the equations, so take 0. 3x0=0. So, with that knowledge, you now only have to solve for m, m must equal 7 then. Haha

jabez177 (jabez177):

The answer is -3.